OpenStudy (anonymous):

A man invests a certain amount of money at 2% interest and $800 more than that amount in another account at 4% interest. At the end of one year, he earned $92 in interest. How much money was invested in each account? $1,500 at 2%; $2,300 at 4% $1,400 at 2%; $2,200 at 4% $1,000 at 2%; $1,800 at 4%

OpenStudy (anonymous):

Let the amount of money invested @ 2% for one year be x. then, Interest2 = Amount*Rate*Time/100 -2x/100 (Amount - x , Rate - 2/100 , Time - 1) Let the amount of money invested @ 4% for one year be (x+800). then, Interest1 = (x+800)4 / 100 Given, total interest recieved = 92 therefore, Interest1 + Interest2 = 92 = 2x/100 + 4(x+800)/100 = 92 2x + 4x + 3200 / 100 = 92 6x = 9200-3200 6x= 6000 x= 1000 therefore, amount 1 = x = $1000 @ 2% and amount 2 = x+800 = $1800 @ 4%

OpenStudy (anonymous):

thx alot